OpenStudy (anonymous):

4(2x -3) + 8(x - 4) = 2(2x + 6)

OpenStudy (jiteshmeghwal9):

\[4(2x-3)+8(x-4)=2(2x+6)\]\[=>8x-12+8x-32=4x+12\]\[=>16x-44=4x+12\] Now solve it & gt ur answer.

OpenStudy (jiteshmeghwal9):

Can u do it from here??

OpenStudy (jiteshmeghwal9):

@Brooke.Summer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8X-12+8X-32=4X+12 16X-44=4X+12 subtract 4X on both sides 12X-44=12 add 44 on both sides 12X=56 divide 12 on both sides X=4.7